# SPRMamba: Surgical Phase Recognition for Endoscopic Submucosal Dissection with Mamba

### v1

@@ 4.5.4 Effect of using convolutions
| Conv without Dilation | 85.9711.48 | 85.7517.70 | 83.9618.06 | 74.7418.92 |
| Without Conv | 86.6010.95 | 85.7117.94 | 84.9917.56 | 75.4618.87 |
| Conv with Dilation | 87.649.83 | 86.7216.54 | 86.7615.66 | 77.5117.83 |
@@ 5 Disscussion
To address these challenges, we propose SPRMamba, a Mamba-based framework for online ESD surgical phase recognition. Our approach leverages the Scale Residual TranMamba module to effectively model both long-term and short-term temporal contexts, offering superior temporal modeling capabilities compared to traditional methods. Additionally, We introduce a novel Temporal Sample Strategy to mitigate the computational resource constraints, making the framework feasible for real-time surgical phase recognition. The advantages of our approach are as follows: 1) Long-Term Modeling Superiority: Mamba enhanced capabilities for long-term temporal modeling with lower computational complexity compared to Transformers. This makes it particularly suitable for handling the extended duration and complexity of ESD surgeries; 2) Effectiveness of the SRTM Module: The proposed SRTM module achieves the best performance due to its ability to capture complex phase relationships in ESD surgeries. While Mamba ensures overall phase recognition accuracy, the challenge of accurately recognizing short-duration phases remains, which our SRTM module addresses effectively; 3) Advantages of the Temporal Sample Strategy The proposed Temporal Sample Strategy outperforms direct long sequence processing and mitigates the secondary complexity typically associated with Transformer models, enhancing the efficiency of temporal context modeling and improving overall recognition performance. In addition, our proposed method allows for flexible modification of the temporal modeling length and specific adjustments for different surgical tasks.
@@
Finally, our proposed method not only surpasses state-of-the-art (SOTA) methods in ESD surgical phase recognition but also demonstrates robustness for other surgical tasks. Specifically, our method achieves an average accuracy of 87.64%, an average precision of 86.72%, an average recall of 86.76%, and an average Jaccard of 77.51% on the ESD dataset, outperforming the next best method by significant margins. Additionally, validation on the cholecystectomy Cholec80 dataset further confirms our method’s robustness, achieving superior results in most metrics compared to the SOTA methods[3, 45, 49, 26, 7, 6, 9, 48], except for precision. The qualitative results, as illustrated in Fig. 3 and Fig. 4, further substantiate the effectiveness and reliability of our approach.
@@
Although our method shows high surgical phase recognition performance in ESD and cholecystectomy videos, it has some limitations. First, this study primarily focuses on ESD procedures, and while we validated the robustness of our method on the Cholec80 dataset, further experiments are needed to assess its generalizability across a broader range of surgical tasks. In the future, we plan to extend our method to other surgical video analysis tasks, potentially incorporating other innovations such as uncertainty analysis to enhance the robustness and accuracy of SPRMamba across different surgical. Second, the dataset used in this study was limited in size and diversity, which may affect the model’s performance in more diverse clinical settings. Future work will focus on expanding the dataset to include multicenter data and a wider variety of surgical procedures.
@@ 6 Conclusion
In this paper, we address the critical need for accurate surgical phase recognition in Endoscopic Submucosal Dissection, a key procedure for treating early-stage gastrointestinal cancers. Achieving precise real-time phase recognition in ESD is essential for improving surgical outcomes and efficiency. However, traditional SPR methods face significant challenges, particularly in effectively capturing the temporal context over extended surgery durations and managing the high computational demands of video analysis required for real-time applications. To overcome these challenges, we propose SPRMamba, a novel framework designed to enhance the accuracy and efficiency of ESD surgical phase recognition. Specifically, our method proposes the Scale Residual TranMamba module, which combines the ability of Mamba to extract the long-term temporal context with the ability of the Transformer to extract the short-term temporal context and excels in capturing complex temporal relationships in surgical videos. In addition, considering the real-time requirements of surgical phase recognition, a temporal sampling strategy is designed to optimize computational resources by efficiently modeling both short-term and long-term temporal contexts. Extensive experiments demonstrate that SPRMamba not only outperforms current state-of-the-art methods in accuracy and robustness but also significantly reduces the computational burden. In the future our approach has the potential to advance the field of surgical video analysis, offering a valuable tool for both clinical practice and surgical education.

### v2

@@ 4.4.6 Effect of Using Convolutions
| Conv | 85.9711.48 | 85.7517.70 | 83.9618.06 | 74.7418.92 |
| w/o Conv | 86.6010.95 | 85.7117.94 | 84.9917.56 | 75.4618.87 |
| Dilation | 87.649.83 | 86.7216.54 | 86.7615.66 | 77.5117.83 |
@@ 5 Disscussion
To address these challenges, we propose SPRMamba, a Mamba-based framework for online ESD surgical phase recognition. Our approach leverages the Scale Residual TranMamba module to effectively model both long-term and short-term temporal contexts, offering superior temporal modeling capabilities compared to traditional methods. Additionally, we introduce a novel Hierarchical Sampling Strategy to mitigate the computational resource constraints, making the framework feasible for real-time surgical phase recognition. Both quantitative and qualitative results form our analytical anlation experiments conducted on the ESD385 dataset demonstrate:
(1) Long-Term Temporal Modeling with Linear Complexity: By integrating Mamba’s state-space architecture, SPRMamba achieves superior long-range dependency modeling (capturing phases up to 20+ minutes) at 27% lower FLOPs than Transformer-based methods. This linear complexity ( vs. ) is critical for real-time ESD analysis, where traditional approaches like MS-TCN [12] and Trans-SVNet [9] fail to balance accuracy and speed (Table 2).
(2) Fine-Grained Phase Transition Resolution via SRTM: The Scale Residual TranMamba (SRTM) module uniquely addresses ESD’s low inter-phase variance (e.g., Incision and ESD phases) by combining Mamba’s global context with Transformer-like local attention. This hybrid design reduces misclassifications in short-duration phases (e.g., Preparation) compared to pure Mamba architectures, as validated in Fig. 8.
(3) Hierarchical Sampling Strategy for Clinical Feasibility: The proposed Hierarchical Sampling Strategy outperforms direct long sequence processing and mitigates the secondary complexity typically associated with Transformer models, enhancing the efficiency of temporal context modeling and improving overall recognition performance. Crucially, it supports adaptive temporal modeling lengths, accommodating variability in ESD workflows across surgeons.
@@
Finally, compared with the SOTA method, our proposed method not only achieves better results in ESD surgical phase recognition, but also demonstrates robustness for other surgical tasks. Specifically, our method achieves an average accuracy of 87.64%, an average precision of 86.72%, an average recall of 86.76%, and an average Jaccard of 77.51% on the ESD dataset, outperforming the next best method by significant margins. Additionally, validation on the cholecystectomy Cholec80 dataset further confirms our method’s robustness, achieving superior results in most metrics compared to the SOTA methods [3, 45, 49, 26, 7, 6, 9, 48], except for precision. The qualitative results, as illustrated in Fig.6 and Fig.7, further substantiate the effectiveness and reliability of our approach.
@@
Although our method shows high surgical phase recognition performance in ESD and cholecystectomy videos, it has some limitations. First, this study primarily focuses on ESD procedures, and while we validated the robustness of our method on the Cholec80 dataset, further experiments are needed to assess its generalizability across a broader range of surgical tasks. In the future, we plan to extend our method to other surgical video analysis tasks, potentially incorporating other innovations such as uncertainty analysis to enhance the robustness and accuracy of SPRMamba across different surgical settings. Second, the dataset used in this study was limited in size and diversity, which may affect the model’s performance in more diverse clinical settings. Future work will focus on expanding the dataset to include multicenter data and a wider variety of surgical procedures.
@@ 6 Conclusion
In this paper, we address the critical need for accurate surgical phase recognition in Endoscopic Submucosal Dissection, a key procedure for treating early-stage gastrointestinal cancers. Achieving precise real-time phase recognition in ESD is essential for improving surgical outcomes and efficiency. However, traditional SPR methods face significant challenges, particularly in effectively capturing the temporal context over extended surgery durations and managing the high computational demands of video analysis required for real-time applications. To overcome these challenges, we propose SPRMamba, a novel framework designed to enhance the accuracy and efficiency of ESD surgical phase recognition. Specifically, our method proposes the Scale Residual TranMamba module, which combines the ability of Mamba to extract the long-term temporal context with the ability of the Transformer to extract the short-term temporal context and excels in capturing complex temporal relationships in surgical videos. In addition, considering the real-time requirements of surgical phase recognition, a temporal sampling strategy is designed to optimize computational resources by efficiently modeling both short-term and long-term temporal contexts. Extensive experiments demonstrate that SPRMamba not only outperforms current state-of-the-art methods in accuracy and robustness but also significantly reduces the computational burden. In the future, our approach has the potential to advance the field of surgical video analysis, offering a valuable tool for both clinical practice and surgical education.